Calculus II Spring 2002 MA114:004–006 Russell Brown

Lecturer: Russell Brown, POT 741, 859 257 3951, rbrown@uky.edu. Office hours: M11-12 in Mathskeller, WF11-12 in POT741 and by appointment. Teaching assistants: Ms. Wendy Martin, sections 004, 006, Office: POT702, 859 257 6804, dmartin@ms.uky.edu, Office hours: TBA. Mr. Daniel Kiteck, section 005, POT802, 859 257 6816, dkiteck@ms.uky.edu, Office hours: TBA. Rooms: The lectures will be in CB234 on MWF. The recitations meet TR as in the schedule book. Textbook: The textbook for this course will be Calculus, 3rd edition, by James Stewart. Material to be covered: The second semester of Calculus will begin with the study of the exponential, logarithm and inverse trigonometric functions. Then, we will study techniques of integration. The third part of the course is devoted to sequences and series. The study of series and their convergence allows us to define new mathematical functions. A good understanding of convergence will help us to understand to the accurracy of numerical approximations. about derivatives, The final part of the course introduces parametric curves and (reviews?) polar coordinates. These topics serve as a preparation to MA213. Homework: Regular homework assignments will be made, but will generally not be collected for marking. Some time will be made available in recitation to go over the problems. The homework and quiz grade will be based on occasional quizzes and homework problems. The list of problems on the syllabus provide guidance as to the topics covered and are suggested for students who feel they need additional practice. Mathematics resource center: Teaching assistants will be in the Mathematics resource center to help with Calculus. This resource center will be in CB065 (in the basement) A schedule will be circulated early in the semester and will eventually be located at http://www.mathskeller.com Exams: There will be three exams and a final. These exams are scheduled in the evening as indicated in the course calendar. Please be sure that you have these dates free. The final exam will be cumulative, but with an emphasis on the material covered since the last test. MA194: In addition, to the 4 hours of credit for MA114, the department offers one additional hour of credit as MA194 on a pass/fail basis. You will pass MA194 if you pass MA114 and have 0, 1 or 2 unexcused absences. If you fail MA114 or have more unexcused absences, you will fail MA194. Your section number for MA194 should equal your section number for MA114. If you drop or change sections of MA114, please make sure to also drop or change sections of MA194. Grading: Students need an average of 90% for an A, 80% for a B, 70% for a C and 60% for a D. Grades may be curved by making small adjustments in these percentages. Your grade will be based on the activities in the table below.

3 hour exams  300
Final exam  100
Homework and quizzes 100
TOTAL  500

Calculators: Students may use a graphing calculator on exams and homework. Several topics such as series, parametric curves and polar coordinates will be much simpler with a graphing calculator. Students may not use a machine with symbolic manipulation capabilities on exams. Thus, no TI-89’s, TI-92’s, no HP-48’s or laptop computers may be used on exams. Please see the lecturer if you have any questions as to whether a particular machine may be used on a test. Absences: You should attend class. If you must miss a recitation and are registered for MA194, you must explain your absence to your teaching assistant. Otherwise, your absence will be marked as unexcused and this may lead to failing MA194. If you are not able to turn in a homework assignment because of an absence, you will not be able to turn it in late. If you have an excused absence which causes you to miss an assignment, please inform your lecturer, Russell Brown. A list of all excused absences will be collected during the semester. We will consult this list before assigning final grades. Web page: A primitive web page for this course is at http://www.ms.uky.edu/∼rbrown/courses/ma114.s.02 Any handouts will be available at this address. Library: Handouts, solutions to exams and occasional homework problems will be kept on reserve in the Mathematics Library in the basement of POT.